# Towards A Universal Graph Structural Encoder

### v1

@@ Abstract
Recent advancements in large-scale pre-training have shown the potential to learn generalizable representations for downstream tasks. In the graph domain, however, capturing and transferring structural information across different graph domains remains challenging, primarily due to the inherent differences in topological patterns across various contexts. Additionally, most existing models struggle to capture the complexity of rich graph structures, leading to inadequate exploration of the embedding space. To address these challenges, we propose GFSE, a universal graph structural encoder designed to capture transferable structural patterns across diverse domains such as molecular graphs, social networks, and citation networks. GFSE is the first cross-domain graph structural encoder pre-trained with multiple self-supervised learning objectives. Built on a Graph Transformer, GFSE incorporates attention mechanisms informed by graph inductive bias, enabling it to encode intricate multi-level and fine-grained topological features. The pre-trained GFSE produces generic and theoretically expressive positional and structural encoding for graphs, which can be seamlessly integrated with various downstream graph feature encoders, including graph neural networks for vectorized features and Large Language Models for text-attributed graphs. Comprehensive experiments on synthetic and real-world datasets demonstrate GFSE’s capability to significantly enhance the model’s performance while requiring substantially less task-specific fine-tuning. Notably, GFSE achieves state-of-the-art performance in $81.6\%$ evaluated cases, spanning diverse graph models and datasets, highlighting its potential as a powerful and versatile encoder for graph-structured data.
Machine Learning, ICML
[…]
1 Introduction
@@
The success of large-scale pre-trained models in domains such as natural language processing (Achiam et al., 2023; Bubeck et al., 2023; Touvron et al., 2023), computer vision (Radford et al., 2021; Ramesh et al., 2021), audio (Yang et al., 2023; Borsos et al., 2023) has demonstrated the power of learning transferable representations across diverse datasets. However, the development of such models remains challenging in the graph domain. Due to the inherent difference in dataset-specific features, most prior graph pre-training models are specialized for certain areas, such as molecules (Zhang et al., 2020; Sypetkowski et al., 2024), proteins (Nijkamp et al., 2023), and knowledge graphs (Galkin et al., 2023). These specialized models require domain-specific knowledge and suffer from limited transferability to different graph domains.
@@
Graphs exhibit rich structural patterns that are often shared across domains. For example, social networks commonly exhibit small-world properties and community structures, while biological networks reveal recurring motifs and hierarchical modularity. Similarly, citation networks share characteristics like scale-free degree distributions and core-periphery structures. Despite their prevalence, they are often overlooked in favor of domain-specific features, limiting the transferability of existing graph pre-trained models.
@@
As tokens in NLP derive meaning from their context, nodes in graphs gain significance through their structural patterns. Therefore, we propose a paradigm shift that focuses on universal structural attributes—domain-agnostic features that are shared across diverse graph domains. These attributes, such as motifs, node degrees, and hierarchical topologies, serve as the building blocks for transferable graph representations. Developing a graph pre-trained model that can capture these universal structural patterns in a foundational manner, while remaining flexible enough to handle domain-specific adaptation, presents significant challenges. These challenges highlight the need for innovative pre-training strategies that prioritize universal graph encoding over domain-specific features.
@@
Proposed work. To address the challenges of cross-domain pre-training and effectively capturing universal structural encoding, we propose GFSE, a Graph Foundational Structural Encoder, as shown in Figure 1. GFSE is pre-trained across diverse graph domains using multiple self-supervised pre-training tasks, including shortest path distance regression, motif counting, local community detection, and graph-level contrastive learning. Each pre-training task targets a critical and necessary aspect of graph structures, enabling GFSE to capture a comprehensive understanding of graph topology. GFSE employs a Graph Transformer enhanced with biased attention mechanisms. Notably, the relative positional encoding, derived from the random walk matrix, is explicitly integrated into the attention bias term. This design allows GFSE to effectively capture intricate structural dependencies among node pairs during pre-training, ensuring both efficiency and theoretically guaranteed expressiveness. GFSE’s versatility extends to various graph learning scenarios. The pre-trained GFSE can produce generic and expressive Positional and Structural Encodings (PSE) for topological tasks. In feature-enriched contexts, the generated PSE can seamlessly augment vectorized features or integrate with text encoders (e.g., LLMs) for text-attributed graphs, enabling GFSE to serve as a powerful component in graph foundation models.
@@
The contributions of this work are threefold. (1) We propose GFSE, the first cross-domain graph structural encoder pre-trained with four essential self-supervised learning objectives. Extensive experiments show the effectiveness of these pre-training tasks, leading to state-of-the-art performance in $81.6\%$ evaluated cases, spanning diverse graph models and datasets. (2) We provide theoretical justification and empirical results demonstrating GFSE’s ability to generate expressive PSE. (3) GFSE serves as a plug-and-play solution for any graph foundation model to incorporate structural information. By focusing on universal structural patterns, GFSE paves the way for more generalizable and adaptable graph encoding, potentially reducing the need for domain-specific fine-tuning in many applications.
@@
Figure 1: A) GFSE is pre-trained on 8 datasets from 6 different domains. Pre-training tasks include reconstruction (shortest path distance regression, motif counting) and contrastive learning (community detection, graph contrastive learning). B) GFSE generates generic and expressive Positional and Structural Encoding (PSE) to tackle topological graph tasks. GFSE can also be seamlessly integrated into downstream feature encoders for feature-enriched tasks by concatenating with initial vectorized features or prepending the generated PSE to the textual prompt as a soft token.
2 Related Work
@@
Graph Pre-training. Graph self-supervised learning approaches are typically pre-training graph models, e.g., GNNs or Graph Transformers, on a massive amount of labeled graphs with inherent features by reconstructing the structures or masked attributes (Cui et al., 2020; Hou et al., 2022; Kipf & Welling, 2016b; Hu et al., 2020b; Wang et al., 2017; Xia et al., 2024; Xia & Huang, 2024; Zhao et al., 2024b; Mizera et al., 2024). Some works also utilize contrastive learning to enhance node and graph-level representation learning (Han et al., 2022; Hassani & Khasahmadi, 2020; Velickovic et al., 2019; Hu et al., 2019; Lee et al., 2022; Li et al., 2021; Lu et al., 2021; Sun et al., 2019, 2021; Xu et al., 2021a; Galkin et al., 2023; Zhao et al., 2024a). These methods, while effective in certain domains, exhibit limited generalizability across different graph domains due to their tailored design for specific types of data. Additionally, there have been graph prompt techniques (Huang et al., 2024; Fang et al., 2024) that can be used to enhance model adaptation over graphs. There have also been some attempts at cross-domain graph pre-training models (Qiu et al., 2020; Davies et al., 2023). Unfortunately, all these models rely on one singular pre-training task (i.e., contrastive learning), and usually fail to capture fine-grained structural features at node level or edge level (Mao et al., 2024).
Graph Language Models. With the success of foundation models in NLP, recent efforts also harness LLMs to develop domain-specific graph foundation models by flattening graph structures and associated texts into prompts (Chen et al., 2024a; Tang et al., 2023; Ye et al., 2023; Qian et al., 2023; Zhao et al., 2023; Guo et al., 2023; Chen et al., 2024b; Liu et al., 2023a; Kong et al., 2024; Chen et al., 2024c; Fan et al., 2024; Zhang et al., 2024; Li et al., 2024). Nevertheless, recent studies show that LLM demonstrates an unsatisfying ability to reason and understand complicated structures within graph (Fatemi et al., 2023; Zhao et al., 2023; Wang et al., 2024). In this work, we focus on developing an expressive graph model dedicated to encoding the rich topological information, without relying on associated text. Our approach complements LLMs on text-attributed graphs, serving as a universal structural encoder.
Positional and Structural Encoding (PSE). Traditional PSEs include hand-crafted features such as Laplacian PE (Davies et al., 2024; Kreuzer et al., 2021; Beaini et al., 2021; Wang et al., 2022), shortest-path distance (Li et al., 2020; Ying et al., 2021), kernel distance (Mialon et al., 2021), random-walk encoding (Ma et al., 2023; Dwivedi et al., 2021; Brüel-Gabrielsson et al., 2022; Rampášek et al., 2022), node degree centrality (Ying et al., 2021), etc. Some studies have introduced specialized networks designed to adaptively learn PSE to enhance performance (Kreuzer et al., 2021; Dwivedi et al., 2021; Chen et al., 2022; Lim et al., 2022). GPSE (Liu et al., 2023b) proposes to pre-train a positional and structural encoder on domain-specific graphs to generate PSE. However, GPSE still suffers from limited transferability and expressiveness across other domains, due to its simplistic pre-training backbone and randomized node features. Finding the most effective and versatile PSE remains an open challenge that requires further innovation.
3 Proposed Method
@@
As shown in Figure 1, we collect graph pre-training datasets from six different domains, including molecules, proteins, social networks, images, product networks, and academic networks. GFSE utilizes a transformer-based architecture with biased attention to incorporate relative inductive bias within graph structures (Sec. 3.1). GFSE is pre-trained with four challenging self-supervision tasks simultaneously, each designed to enhance a crucial aspect of structural awareness and promote encoding quality (Sec. 3.2). GFSE generates expressive positional and structural encoding (PSE) for topological tasks. The generated PSE can also be seamlessly integrated into graphs with vectorized features or textual features, enhancing the downstream performance (Sec. 3.3).
3.1 Architecture
@@
Previous work (Liu et al., 2023b) uses randomized features to replace initial node features. However, it leads to poor generalizability across different domains. In this work, we propose to use both absolute and relative random-walk positional encoding as the initial features. Formally, let $G(V,E)$ represent an input graph, where $V$ and $E$ denote the set of nodes and edges, respectively. $\mathbf{A}\in\mathbb{R}^{N\times N}$ indicates the adjacency matrix, where $N$ is the number of nodes, and $\mathbf{D}$ is the degree matrix. Random Walk matrix is defined as $\mathbf{M}=\mathbf{D}^{-1}\mathbf{A}$, where $\mathbf{M}_{i,j}$ indicates the transition probability from the $i$-th node to the $j$-th node. Following previous works (Ma et al., 2023), we calculate the $d$-dimensional encoding for each node and node pairs as follows.
$\mathbf{P}_{i}$ $=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\cdots,\mathbf{M}^{d-1}]_{i,i}$ (1)
$\mathbf{R}_{i,j}$ $=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\cdots,\mathbf{M}^{d-1}]_{i,j}$ (2)
@@
Pre-training Backbone. GFSE is built on a GPS architecture (Rampášek et al., 2022) for pre-training, due to its scalability and generalizability. Each GPS layer contains local message passing and global attention modules to capture neighbor and long-range information. In the $\ell$-th layer, the node encoding $\mathbf{P}^{\ell}$ and relative edge encoding $\mathbf{R}^{\ell}$ are fed into message passing layers ($\operatorname{MPNN}$) and Biased Attention Module ($\operatorname{BiasAttn}$) parallelly. The node encoding $\mathbf{P}^{\ell+1}$ is updated by aggregating local and global information.
$\vspace{-0.5cm}\mathbf{P}_{M}^{\ell+1},\mathbf{R}^{\ell+1}$ $=\operatorname{MPNN}^{\ell}\left(\mathbf{P}^{\ell},\mathbf{R}^{\ell},\mathbf{A}\right)$ (3)
$\mathbf{P}_{T}^{\ell+1}$ $=\operatorname{BiasAttn}^{\ell}\left(\mathbf{P}^{\ell},\mathbf{R}^{\ell}\right)$ (4)
$\mathbf{P}^{\ell+1}$ $=\operatorname{MLP}^{\ell}\left(\mathbf{P}_{M}^{\ell+1}+\mathbf{P}_{T}^{\ell+1}\right)$ (5)
@@
Attention Bias. The global attention in the original GPS framework does not account for relative edge encoding while leaving them entirely for the message-passing layers. However, incorporating relative edge encoding in global attention is crucial for capturing long-range dependencies, as the receptive field of message-passing layers is inherently constrained by their depth. To build a theoretically more powerful GPS, we explicitly incorporate relative edge encoding into global attention, where the attention weight between the $i$-th and the $j$-th nodes is computed by $a_{i,j}^{\prime}=\operatorname{SoftMax}(a_{i,j}+\operatorname{Linear}(\mathbf{R}_{i,j}))$, where $\operatorname{Linear}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ indicates a linear layer that maps the $d$-dimensional relative encoding to a scalar. $a_{i,j}$ denotes the original attention weight computed by scaled-dot self-attention on the node encoding $\mathbf{P}^{\ell}$ in each GPS layer. We refer to Appendix C.3 (Figure 5) for biased attention details.
3.2 Self-supervised Pre-training Tasks
GFSE is pre-trained with reconstruction and contrastive structural tasks at multiple context levels. Each task highlights a specific structural aspect, thereby augmenting the model’s expressiveness and capability to capture complex graph structures. Let $\mathbf{P}^{L}\in\mathbb{R}^{N\times d_{e}}$ represent the output after $L$ GPS layers. We decode $\mathbf{P}^{L}$ with independent MLP heads for each pre-training task.
[…]
Motif Counting enables the model to better identify each node’s role in the surrounding subgraphs. To improve expressiveness, we include a variety of small motifs, called graphlets (Pržulj et al., 2004; Pržulj, 2007; Bouritsas et al., 2022), with different numbers of vertices, which are beyond usual types like stars, paths, cycles, and cliques. We refer to Appendix B.2 for more technical details. Let $Y_{i}^{m}\in\mathbb{Z}^{k}$ denote the motif label that indicates the number of certain motifs surrounding node $i$, where $k$ is the number of graphlet types. The loss is formulated as $\mathcal{L}_{\text{MC}}=\frac{1}{|V|}\sum_{i\in V}\|h_{\text{MC}}(\mathbf{P}^{L}_{i})-Y_{i}^{m}\|^{2}$, where $h_{\text{MC}}:\mathbb{R}^{d}_{e}\rightarrow\mathbb{R}^{k}$ is the task-specific head for motif counting.
Community Detection aims to identify densely connected subgraphs, where nodes within a community are more closely linked to each other than to nodes outside the community. Such community structures are ubiquitous in various real-world networks, e.g., social networks, and transportation systems. We employ the Louvain Community Detection Algorithm (Blondel et al., 2008) to extract the community structure from pre-training graphs, which clusters nodes into communities based solely on graph topology without node features. We approach this task in a contrastive learning manner by minimizing the embedding distances between intra-community nodes while maximizing the distance between inter-community nodes.
$\mathcal{L}_{\text{CD}}=$ $\sum_{i\in V}\sum_{j\in V}Y_{i,j}^{c}(1-\operatorname{sim}(i,j))$
$+(1-Y_{i,j}^{c})\max(0,\epsilon-(1-\operatorname{sim}(i,j))),$
where the similarity score $\operatorname{sim}(i,j)$ is calculated by $\operatorname{sim}(i,j)=\frac{\boldsymbol{z}_{i}\cdot\boldsymbol{z}_{j}}{||\boldsymbol{z}_{i}||\cdot||\boldsymbol{z}_{j}||}$ and $\boldsymbol{z}_{i}=h_{\text{CD}}(\mathbf{P}^{L}_{i})$ with a head $h_{\text{CD}}$. $\epsilon$ is a margin hyperparameter. $Y_{i,j}^{c}$ is a binary label that indicates if the $i$-th node and the $j$-th node are in the same community. Through $\mathcal{L}_{\text{CD}}$, GFSE learns to discern community boundaries and distinguish nodes based on local community structures.
Graph Contrastive Learning aims to distinguish graphs from different datasets. The motivation is rooted in the observation that similar structural characteristics in different domains may exhibit distinct meanings. For example, a subgraph representing a protein interaction network might correspond to a specific functional module, whereas a similar structure in a social network could represent a tightly-knit community or interest group. Therefore, GFSE distinguishes graphs from different datasets, in addition to performing structural pre-training tasks within a single graph. The loss for this task is computed by
$\mathcal{L}_{\text{GCL}}=-\log\frac{\exp\left(\operatorname{sim}\left(\boldsymbol{z}_{G_{i}},\boldsymbol{z}_{G_{j}}\right)/\tau\right)}{\sum_{k=1}^{K}\mathbbm{1}_{[G_{k}\nsim G_{i}]}\exp\left(\operatorname{sim}\left(\boldsymbol{z}_{G_{i}},\boldsymbol{z}_{G_{k}}\right)/\tau\right)},$
where $\tau$ is the temperature, $G_{i}$ and $G_{j}$ are from the same dataset, $\boldsymbol{z}_{G_{i}}=\mathrm{GlobalPool}(h_{\text{GCL}}(\mathbf{P}^{L}_{G_{i}}))$ is the output of the global pooling applied to the final layer’s representation $\mathbf{P}^{L}_{G_{i}}$ for the graph $G_{i}$, $K$ is the number of negative samples, and $\mathbbm{1}_{[G_{k}\nsim G_{i}]}$ is an indicator function that determines whether graphs $G_{k}$ and $G_{i}$ originate from different datasets.
Multi-task Selection and Loss Balance. Each pre-training task targets a different structural aspect, enabling GFSE to capture a comprehensive understanding of graph topology. For instance, the shortest path distance regression task emphasizes global connectivity within graphs, while motif counting focuses on the occurrence of local subgraph patterns. Although other pre-training tasks could be considered, the chosen tasks provide a balanced and effective combination, capable of capturing both local and global structural information at varying granularities without introducing unnecessary complexity. Since the loss scale and difficulty vary significantly across tasks, we introduce task-specific uncertainty (Kendall et al., 2018), which is learnable during pre-training to unify the scales of all losses. Task-specific uncertainty is used to automatically balance different pre-training losses, i.e., $\mathcal{L}_{\text{SPD}},\mathcal{L}_{\text{MC}},\mathcal{L}_{\text{CD}}$, and $\mathcal{L}_{\text{GCL}}$ (see Appendix C.4 for more details). Moreover, the evolution of uncertainty values provides insights into each task’s contribution to the overall pre-training process.
3.3 Combination with Downstream Feature Encoder
Application on Text-attributed Graphs. Language models are typically employed to process the text-attributed graphs, where GFSE can be seamlessly applied to incorporate structural information within graphs. Given the generated $\mathbf{P}^{L}\in\mathbb{R}^{N\times d_{e}}$, an MLP is employed to project $\mathbf{P}^{L}$ into the embedding space of the language model. This projected PSE is then prepended as a soft token to the associated text, effectively incorporating the graph’s structural information into the model input. Subsequently, these structure-enriched tokens are fed into downstream large language models (LLMs), enhancing their performance on graph-related tasks. The process, involving training a lightweight MLP and fine-tuning LLM with Parameter-Efficient Fine-Tuning (PEFT) techniques such as LoRA (Hu et al., 2021), makes it scalable and efficient for large-scale text-attributed graph applications.
3.4 Expressive Power of GFSE

### v2

@@ Abstract.
Recent advancements in large-scale pre-training have shown the potential to learn generalizable representations for downstream tasks. In the graph domain, however, capturing and transferring structural information across different graph domains remains challenging, primarily due to the inherent differences in graph topological patterns across various contexts. For example, a social network’s structure is fundamentally different from that of a product co-purchase graph. Additionally, most existing models struggle to capture the rich topological complexity of graph structures, leading to inadequate exploration of the graph embedding space. To address these challenges, we propose GFSE, a universal pre-trained graph encoder designed to capture transferable structural patterns across diverse domains such as the web graph, social networks, and citation networks. GFSE is the first cross-domain graph structural encoder pre-trained with multiple self-supervised learning objectives. Built on a Graph Transformer, GFSE incorporates attention mechanisms informed by graph structural information, enabling it to encode intricate multi-level and fine-grained topological features within complex graph structures. The pre-trained GFSE produces generic and theoretically expressive positional and structural encoding for graphs, which can be seamlessly integrated with various downstream graph feature encoders, including graph neural networks for vectorized features and Large Language Models (LLMs) for text-attributed graphs. Comprehensive experiments on synthetic and real-world datasets demonstrate GFSE’s capability to significantly enhance the model’s performance while requiring substantially less task-specific fine-tuning. Notably, GFSE achieves state-of-the-art performance in diverse evaluated cases spanning various graph models and datasets, highlighting its potential as a powerful and versatile encoder for graph-structured data.
Graph Representation Learning, Self-supervised Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792656isbn: 979-8-4007-2307-0/2026/04ccs: Computing methodologies Supervised learning
1. Introduction
@@
The success of large-scale pre-trained models in domains such as natural language processing (achiam2023gpt; bubeck2023sparks; touvron2023llama), computer vision (radford2021learning; ramesh2021zero), audio (yang2023uniaudio; borsos2023audiolm) has demonstrated the power of learning transferable representations across diverse datasets. However, developing such foundational models for the graph domain remains a significant challenge. Due to the inherent heterogeneity of graph-structured data across the domains, most prior graph pre-training efforts are highly specialized, focusing on narrow domains such as link structures in web graphs, user interactions in social networks, or relational facts in knowledge graphs (galkin2023towards; zhang2025litfm; xia2024anygraph). These specialized models often rely heavily on domain-specific features, which limits their transferability and requires deep domain expertise to adapt.
@@
This feature-centric approach overlooks a critical, unifying aspect of graphs: they exhibit rich structural patterns that are often shared across domains. For example, social networks and citation networks commonly exhibit small-world properties and dense community structures, while e-commerce graphs reveal hierarchical taxonomies. Despite the prevalence of these fundamental patterns, they are often secondary to domain-specific features, hindering the generalization capabilities of existing pre-trained models.
@@
As tokens in NLP derive meaning from their context, nodes in graphs gain significance through their structural role. Therefore, we propose a paradigm shift that prioritizes learning universal structural attributes—domain-agnostic features that are shared across diverse graph domains. These attributes, such as motifs, node degrees, and hierarchical topologies, serve as the building blocks for transferable graph representations. The central challenge lies in developing a foundational model that can effectively capture these universal structural patterns while remaining adaptable for downstream tasks. Such a model would reduce the reliance on domain-specific engineering and unlock the potential for truly general-purpose graph representation learning.
@@
Proposed work. To address the challenges of cross-domain pre-training and effectively capturing universal structural encoding, we propose GFSE, a Graph Foundational Structural Encoder, as shown in Figure 1. GFSE is pre-trained across diverse graph domains using multiple self-supervised pre-training tasks, including shortest path distance regression, motif counting, local community detection, and graph-level contrastive learning. Each pre-training task targets a critical and necessary aspect of graph structures, enabling GFSE to capture a comprehensive understanding of graph topology. GFSE employs a Graph Transformer enhanced with biased attention mechanisms. Notably, the relative positional encoding, derived from the random walk matrix, is explicitly integrated into the attention bias term. This design allows GFSE to effectively capture intricate structural dependencies among node pairs during pre-training, ensuring both efficiency and theoretically guaranteed expressiveness. GFSE’s versatility extends to various graph learning scenarios. The pre-trained GFSE can produce generic and expressive Positional and Structural Encodings (PSE) for topological tasks. In feature-enriched contexts, the generated PSE can seamlessly augment vectorized features or integrate with language models for text-attributed graphs, enabling GFSE to serve as a powerful component in graph foundation models.
@@
The contributions of this work are threefold. (1) We propose GFSE, a unified cross-domain graph structural encoder pre-trained with four essential self-supervised learning objectives. Extensive experiments demonstrate the effectiveness of these pre-training tasks, resulting in state-of-the-art performance across diverse evaluated cases, encompassing various graph models and datasets. (2) We provide theoretical justification and empirical results demonstrating GFSE’s ability to generate expressive PSE. (3) GFSE serves as a plug-and-play solution for any graph model to incorporate generalizable structural information, reducing the need for domain-specific fine-tuning on heterogeneous web-scale tasks.
@@
Figure 1. A) GFSE is pre-trained on 8 datasets from 6 different domains. Pre-training tasks include reconstruction and contrastive learning. B) GFSE generates generic and expressive Positional and Structural Encoding (PSE) to tackle topological graph tasks, which can also be seamlessly integrated into downstream feature encoders for feature-enriched tasks by concatenating with initial vectorized features or prepending the generated PSE to the textual prompt as a soft token.
2. Related Work
@@
Graph Pre-training. Graph self-supervised learning approaches are typically pre-training graph models, e.g., GNNs or Graph Transformers, on a massive amount of labeled graphs with inherent features by reconstructing the structures or masked attributes (cui2020adaptive; hou2022graphmae; kipf2016variational; hu2020gpt; wang2017mgae; xia2024opengraph; xia2024anygraph; zhao2024graphany; mizera2024graph). Some works also utilize contrastive learning to enhance node and graph-level representation learning (han2022g; hassani2020contrastive; velickovic2019deep; hu2019strategies; lee2022augmentation; li2021pairwise; lu2021learning; sun2019infograph; sun2021mocl; xu2021infogcl; galkin2023towards; zhao2024all). These methods, while effective in certain domains, exhibit limited generalizability across different graph domains due to their tailored design for specific types of data. Additionally, there have been graph prompt techniques (huang2024prodigy; fang2024universal) that can be used to enhance model adaptation over graphs. There have also been some attempts at cross-domain graph pre-training models (qiu2020gcc; davies2023its). Unfortunately, all these models rely on one singular pre-training task (i.e., contrastive learning), and usually fail to capture fine-grained structural features at the node level or edge level (mao2024graph).
Generalizable Positional and Structural Encoding (PSE). Early research on the web graph focused on characterizing its macroscopic properties, such as power-law degree distributions and the prevalent ’bow-tie’ structure (broder2000graph; faloutsos1999power; albert1999internet), which led to seminal node-importance algorithms like PageRank (brin1998anatomy). This perspective inspired a family of traditional, hand-crafted Positional and Structural Encodings (PSEs) based on such structural statistics, including Laplacian PE (davies2024generalised; kreuzer2021rethinking; beaini2021directional; wang2022equivariant), shortest-path distance (li2020distance; ying2021transformers), random-walk encodings (ma2023graph; dwivedi2021graph; bruel2022rewiring; rampavsek2022recipe), and node degree (ying2021transformers). However, these global metrics and engineered features often fail to capture the complex, local connectivity patterns and subgraph motifs that signify sophisticated behaviors like spam link farms or coordinated user activity (milo2002network; leskovec2007graph). To overcome these limitations, recent work has shifted towards models that learn PSE adaptively (kreuzer2021rethinking; dwivedi2021graph; chen2022structure; lim2022sign). While some methods like GPSE (liu2023graph) pre-train a structural encoder, they still suffer from limited cross-domain transferability due to simplistic backbones. Developing an effective and versatile PSE remains an open challenge that requires further innovation.
Graph Language Models. With the success of foundation models in NLP, recent efforts also harness Large Language Models (LLMs) to develop domain-specific graph foundation models. This typically involves converting graph data into a text-compatible format by linearizing the graph’s structure (e.g., as a sequence of adjacency lists or edge sets) and verbalizing node/edge attributes to create descriptive prompts (chen2024llaga; tang2023graphgpt; ye2023natural; qian2023can; zhao2023graphtext; guo2023gpt4graph; chen2024exploring; liu2023one; kong2024gofa; chen2024text; fan2024graph; zhang2024graphtranslator; li2024graph). The goal is to create a single, powerful model that can reason over both the semantics of text attributes and the graph’s topology. However, this approach faces a fundamental challenge: LLMs, which are pre-trained on sequential text, struggle to comprehend the complex, non-Euclidean nature of graph structures. Recent empirical studies and benchmarks consistently show that serializing a graph into a one-dimensional sequence leads to a significant loss of topological information, impairing an LLM’s ability to reason about complex connectivity, distances, and higher-order patterns (fatemi2023talk; wang2024can; wang2024nlgraph). Our work addresses this gap directly. Instead of relying on a single model for both text and structure, we focus on developing an expressive graph model dedicated solely to encoding rich topological information. Our approach is designed to be complementary to LLMs, serving as a universal structural encoder that can provide the crucial topological context that LLMs inherently lack, thereby enabling a more powerful and synergistic approach for text-attributed graphs.
3. Proposed Method
@@
As shown in Figure 1, we collect graph pre-training datasets from six different domains, including social networks, product networks, academic networks, images, molecules, and proteins. GFSE utilizes a transformer-based architecture with biased attention to incorporate relative inductive bias within graph structures (Sec. 3.1). GFSE is pre-trained with four challenging self-supervision tasks simultaneously, each designed to enhance a crucial aspect of structural awareness and promote encoding quality (Sec. 3.3). GFSE generates expressive positional and structural encoding (PSE) for topological tasks (Sec. 3.2). The generated PSE can also be seamlessly integrated into graphs with vectorized features or textual features, enhancing the downstream performance (Sec. 3.4).
3.1. Architecture
@@
Previous work (liu2023graph) uses randomized features to replace initial node features. However, it leads to poor generalizability across different domains. In this work, we propose to use both absolute and relative random-walk positional encoding as the initial features. Formally, let $G(V,E)$ represent an input graph, where $V$ and $E$ denote the set of nodes and edges, respectively. $\mathbf{A}\in\mathbb{R}^{N\times N}$ indicates the adjacency matrix, where $N$ is the number of nodes, and $\mathbf{D}$ is the degree matrix. Random Walk matrix is defined as $\mathbf{M}=\mathbf{D}^{-1}\mathbf{A}$, where $\mathbf{M}_{i,j}$ indicates the transition probability from the $i$-th node to the $j$-th node. Following previous works (ma2023graph), we calculate the $d$-dimensional encoding for each node and node pairs as follows.
(1) $\mathbf{P}_{i}=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\cdots,\mathbf{M}^{d-1}]_{i,i},\quad\mathbf{R}_{i,j}=[\mathbf{I},\mathbf{M},\mathbf{M}^{2},\cdots,\mathbf{M}^{d-1}]_{i,j}$
@@
Pre-training Backbone. GFSE is built on a GPS architecture (rampavsek2022recipe) for pre-training due to its scalability and generalizability. Each GPS layer contains local message passing and global attention modules to capture neighbor and long-range information. In the $\ell$-th layer, the node encoding $\mathbf{P}^{\ell}$ and relative edge encoding $\mathbf{R}^{\ell}$ are fed into message passing layers ($\operatorname{MPNN}$) and Biased Attention Module ($\operatorname{BiasAttn}$) parallelly. The node encoding is updated by aggregating local and global information $\mathbf{P}^{\ell+1}=\operatorname{MLP}^{\ell}\left(\mathbf{P}_{M}^{\ell+1}+\mathbf{P}_{T}^{\ell+1}\right)$ where
(2) $\mathbf{P}_{M}^{\ell+1},\mathbf{R}^{\ell+1}=\operatorname{MPNN}^{\ell}\left(\mathbf{P}^{\ell},\mathbf{R}^{\ell},\mathbf{A}\right);\quad\mathbf{P}_{T}^{\ell+1}=\operatorname{BiasAttn}^{\ell}\left(\mathbf{P}^{\ell},\mathbf{R}^{\ell}\right).$
@@
Attention Bias. The global attention in the original GPS framework does not account for relative edge encoding while leaving them entirely for the message-passing layers. However, incorporating relative edge encoding in global attention is crucial for capturing long-range dependencies, as the receptive field of message-passing layers is inherently constrained by their depth. To build a theoretically more powerful encoder, we explicitly incorporate relative edge encoding into global attention, where the attention weight between the $i$-th and the $j$-th nodes is computed by $a_{i,j}^{\prime}=\operatorname{SoftMax}(a_{i,j}+\operatorname{Linear}(\mathbf{R}_{i,j}))$, where $\operatorname{Linear}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ indicates a linear layer that maps the $d$-dimensional relative encoding to a scalar. $a_{i,j}$ denotes the original attention weight computed by scaled-dot self-attention on the node encoding $\mathbf{P}^{\ell}$ in each GPS layer.
3.2. Expressive Power of GFSE